Consider equally risky, all-equity financed firms G and D. Both firms are currently trading at $50 per share. Firm G pays no dividends, but firm D pays all its earnings as dividends.
In a year from today, the stock of firm G is expected to be at $65 per share. Firm D is expected to pay a $15 dividend per share at the year-end and its stock price at the year-end is expected to be $50 per share. Capital gains are taxed at half the 40% regular income tax rate, but the dividend income is taxed at 30%.


Required
At what share price should the stock of firm D be trading to-day in a competitive stock market? Show calculations to support your answer.
Answer #1

Both the cos are similar in term to return given, the only difference is that one co gives out Dividend but other gives the Capital Gain after retaining all the Dividend and reinvesting it.

Firm D Dividend after tax = 15 (1-0.30) = 10.5

Firm G Capital Gain after tax = 15 * (1-0.40/2) = 12

Return given by the Firm D = After tax Capital Gain / Price

= 12 / 50

= 24%

Now, Similar returns should be given by the Firm G.

Therefore,

Return = After tax Dividend / Price

24% = 10.5 / Price

Price = 10.5 / 0.24

Price =$43.75

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.